Fermi National Laboratory
 

Operation Summary for November 2, 2009 - November 4, 2009

 

Monday Nov. 2

The day shift began with Operations monitoring the Tevatron (TeV) store 7317 (48.46E30) and the Recycler stash (155.4E10), with the Antiproton Source stacking (76.2E10), with the Linac, Booster, and Main Injector (MI) operating normally, with the NuMI and MiniBooNE experiments taking beam, and with the Collider Detector at Fermilab (CDF) and D-Zero (D0) experiments using the colliding beams.

During the day shift, Operations terminated store 7317 at 12:49 PM. Operations put the TeV into shot setup and established store 7318 at 2:50 PM with an average initial luminosity (AIL) of 109E30. Pbar stacked antiprotons all shift and stashed to the Recycler during beamline optics studies at the end of the shift. Recycler personnel conducted emittance studies from 2:45 to 3:30 PM. The Booster chopper (B:CHOP) caused some problems with beam into Booster; Booster personnel changed out some resistors to fix the problem. MTA - MuCool and Operations personnel conducted timeline and beam-switch studies. Recycler personnel completed ECool repairs to the Pelletron.

During the evening shift, a H- Source vacuum pump tripped off; it reset. Two Linac RF stations (LRF2 and LRF3) tripped off. They reset.

Tuesday Nov. 3

The midnight shift began with Operations monitoring store 7318 (42.45E30) and the stash (292.6E10), with Pbar stacking (28.5E10), and with NuMI and MiniBooNE taking beam.

During the midnight shift, Operations terminated store 7318 at 7:30 AM. Cryo system technicians worked on the TeV sector A2 high energy leads. Operations put the TeV into shot setup after the Cryo techs completed their work.

During the day shift, the TeV shot setup was put on hold in order to ramp the TeV to check the new lead flow valve at sector A2. Operations established store 7320 at 10:30 AM with an AIL of 250E30. MI personnel conducted slipstacking studies. There was some downtime due to water skid work on the H- Source, work on MI power supply ducts, as well as a stuck Klixon on a MI RF station (MIRF18).

During the evening shift, Operations terminated store 7320 at 8:33 PM and then put the TeV into shot setup. Operations established store 7321 at 10:53 PM with an AIL of 228.77E30. Recycler personnel began a beam diffusion study.

Wednesday Nov. 4

The midnight shift began with Operations monitoring store 7321 (170.03E30) and the stash (9.9E10), with Pbar stacking (85.8E10), and with NuMI and MiniBooNE taking beam.

During the midnight shift, Operations reported some downtime due to a Booster East Anode Power Supply trip on an AC input overload fault; it reset.

The Plans for Wednesday

The plans for today are to stack, stash, and store. The Recycler has been having beam emittance problems possibly due to poor vacuum. The Run Coordinator will allow a MI access after the next store to investigate. Booster and Linac will take advantage of this to make some short accesses into their beam enclosures for necessary maintenance work.

 

Accelerator Update Archive

More Information
For Tevatron luminosity charts and the current status of Fermilab's accelerators and detectors (live!), please go to Fermilab Now.

Comments and Suggestions
What do you think about the Accelerator Updates? Please send comments and suggestions to: accelupdates@fnal.gov.

last modified 11/04/2009    email Fermilab
Security, Privacy, LegalFermi National Accelerator Laboratory